A large national bank charges local companies for using their services. A bank official reported the results of a regression analysis designed to predict the bank's charges (Y)—measured in dollars per month—for services rendered to local companies. One independent variable used to predict service charges to a company is the company's sales revenue (X)—measured in millions of dollars. The results of the simple linear regression are provided below. Y = 2,700 + 20X Based on this model, if a company’s sales revenue is 60,000,000$, how much is predicted be charged for using their service?

A. $60,000,000

B. $3900

C. $600

D. $2700
Answer #1

Answer

Given regression equation is y = 2700 +20(x)

we have to find the value of y when x = 60(in millions)

Setting x = 60, we get

y = 2700 + (20*60)

= 2700 + 1200

= $3900

option B is correct